A teacher wants to know whether the format of the assigned textbook affects grades. The teacher randomly assigns 6 students to a paper (hard copy) version of the textbook, 6 students to an online textbook, and 6 students to a control condition with no textbook. The teacher gives a quiz on one of the textbook chapters and obtains the following data. Use this data to answer the following questions 5. Paper Textbook 15 10 12 No TexibookG 21 N 18 12 13 10 X2 2515 10 12 T2 12 12 10 14 a) Compute T1, T2, T3, G, N, & k (6x.5 3 points) M) Compute ss,otal, ss,etween and SSwithin (3 points) SS betwen 12 2112 t0+ 1 8 c) Compute df total, d f within and df between (3 points) d Compute MSwithin and MS between (2 points) DThis is coreet sien yaur Compute the F-statistic (1 point) answer to Sb
Answer #1

(a) As calculated G = 211/18 = 11.72, M1 = T1/6 = 12.5, M2 = T2/6 = 12 and M3 = T3/6 = 10.667

k = 3, N = 18. SS total = 41.6

(b) SS Between = SUM[ n *(mean - overall mean)2] = 6 * (12.5 - 11.72)2 + 6 * (12 - 11.72)2 + 6 * (10.667 - 11.72)2 = 3.6286 + 0.4626 + 6.6824 = 10.7736 \approx     10.77

Therefore SS within = SStotal - SS between = 41.6 - 10.77 = 30.83

(c) dfbetween = k - 1 = 3 - 1 = 2

df within = N - k = 18 - 3 = 15

df total = dfbetween + dfwithin = 2 + 15 = 17

(d) MS between = SS between/dfbetween = 10.77/2 = 5.39

MSwithin = SSwithin/dfwithin = 30.83/15 = 2.06

(e) Therefore F = MS between/ MS within = 5.39/2.06 = 2.62